Please help me solve this equation. 13x-5y=8
3x+5y=8

Answer by ankor@dixie-net.com(22740)

Solving this system of equations:
13x - 5y = 8
3x + 5y = 8
--------------adding eliminates y, leaving x to solve:
16x + 0y = 16
16x = 16
x = 16/16
x = 1
:
Find y by substituting 1 for x in the 2nd equation:
3(1) + 5y = 8
3 + 5y = 8
5y = 8 - 3
5y = 5
y = 5/5
y = 1